OpenStudy (anonymous):

0.4(x-2)=6.0 can you show me steps

OpenStudy (saifoo.khan):

\[0.4x - 0.4(2) = 6.0\] \[0.4x - 0.8 = 6.0\] \[0.4x = 6 + 0.8\] \[0.4x = 6.8\] \[4x = 68\]

OpenStudy (anonymous):

0.6(x-2)=12.0 can you do these steps

OpenStudy (saifoo.khan):